Trading in the 90-day bank-accepted bill contract

On Tuesday 3 April 2018, you will be required to enter into four June 2018 90-day bank accepted bill contracts. You may enter into these contracts as a buyer or as a seller. Whether you enter into these contracts as a buyer or a seller will depend on your expectations as to the likely direction of short-term interest rates. Again, you should state a logical basis for entering into these contracts as a buyer or seller. For example, you might speculate that short-term interest rates are likely to rise or fall. Note that again the basis of your speculation is of lesser importance here than is demonstrating that you understand fully the nature of the transactions that you enter into. Using the closing price of the June 2018 90-day bank accepted bill contract on Tuesday 3 April 2018, detail your financial position after you have entered into these four contracts. You should ignore margin calls but you should include in your discussion the deposit that you were required to provide. On Monday 7 May 2018, you must close out your position and do so at the closing price of the June 2018 90-day bank accepted bill contract. Provide a full report of the profit or loss that you made and explain why this occurred.

Note: opening price on 03/04/18 according to ASX is 98.100

closing price on 03/04/18 according to ASX is 98.071

on 07/05/18 closing price is 98.040
Answer #1

whether to buy or sale contract depend upon the expectation of change in interest rate. If the interest rate is going to rise in future than we should sale contract so that we have to pay less interest than what was prevailing in market in fututre.

Similarly, if the interest is going to fall in future than we should buy contract .So, that we can earn more interest than the rate prevailing in future

Thus the decision of whether to buy or sale contract depend upon the expectation of rise or fall in interest rate

In the given case price on 03/04/18 is 98.100 and if there is expectation of fall in price than we should sale the contract(short).As closing price is less we should sale the contract

financial position for four contract is short.i.e short 4*98.1=392.4

On 7th may when the position is closed:

profit=(98.1-98.040)*4=0.24

profit of 0.24 is made.As we short the contract when the price was 98.100 .So, when the price fall we will gain and when price rise we will be at loss.

As price fall from 98.100 to 98.040 we gain 0.06 on each contract. Hence shorting contract is profitable